OpenStudy (anonymous):

pepperidge farm goldfish is a snack food for which 40% of its calories come from fat. Rold Gold pretzels receive 9% of their calories from fat. how many grams of each would be needed to make 620g of a snack mix for which 15% of the calories are from fat?

OpenStudy (anonymous):

Let g be the grams of goldfish. Let p be grams of pretzels. It says g + p = 620 It also says 0.4*g + 0.09*p = 0.15*620 Solve for g and p.